# Test Time Learning for Time Series Forecasting

### v1

@@ 5.4 Results and Statistical Comparisons for Increased Prediction and Sequence Lengths
Table 5: Average MSE and MAE for different prediction lengths comparing TimeMachine and TTT architectures.
@@
Both shorter and longer sequence lengths affect model performance differently. Shorter sequence lengths (e.g., 2880) provide better accuracy for shorter prediction horizons, with the TTT model achieving an MSE of 0.332 and MAE of 0.356 at a 192-step horizon, outperforming TimeMachine. Longer sequence lengths (e.g., 5760) result in higher errors, particularly for shorter horizons, but TTT remains more resilient, showing improved performance over TimeMachine. For shorter prediction lengths (96 and 192), TTT consistently yields lower MSE and MAE compared to TimeMachine. As prediction lengths grow to 720, both models experience increasing error rates, but TTT maintains a consistent advantage. For instance, at a 720-step horizon, TTT records an MSE of 0.517 compared to TimeMachine’s 0.535. Overall, TTT consistently outperforms TimeMachine across most prediction horizons, particularly for shorter sequences and smaller prediction windows. As the prediction length increases, TTT’s ability to manage long-term dependencies becomes increasingly evident, with models like Conv Stack 5 showing stronger performance at longer horizons.
@@ 5.5 Evaluation
  • Impact of Hidden Layer Architectures: While stacked convolutional architectures, such as Conv Stack 3 and Conv Stack 5, provided incremental improvements, simpler architectures like Conv 3 and Conv 5 also delivered competitive performance. Conv Stack 5 showed a reduction in MSE compared to TimeMachine, at horizon 96, where it achieved an MSE of 0.261 versus TimeMachine’s 0.262.
@@
In summary, our experiments reveal that the TimeMachine-TTT model, especially with architectures like Conv Stack 5 and Conv 3, significantly outperforms existing models when larger sequence and prediction lengths are employed. These findings highlight the importance of increasing both sequence and prediction lengths in time series forecasting, particularly when utilizing architectures capable of dynamically capturing long-term dependencies. Even modest modifications, such as Conv 3, can lead to substantial improvements, emphasizing the effectiveness of simpler yet powerful architectures.
@@ 6 Conclusion and Future Work
In this work, we improved the state-of-the-art (SOTA) model for time series forecasting by replacing the Mamba modules in the original TimeMachine model with Test-Time Training (TTT) modules, which use linear RNNs to capture long-range dependencies. Extensive experiments showed that the TTT architectures—MLP and Linear—performed well, with MLP slightly outperforming Linear. Exploring alternative architectures, particularly Conv Stack 5, improved performance at longer prediction horizons. The most significant gains came from increasing sequence and prediction lengths, where our TTT models consistently matched or outperformed the SOTA model, particularly on larger datasets like Electricity, Traffic, and Weather, highlighting the model’s strength in handling long-range dependencies. While convolutional stacks showed promise, further improvement is possible by refining hidden layer configurations and exploring architectural diversity. Overall, this work demonstrates the potential of TTT modules in long-term forecasting, especially with larger datasets and longer horizons.

### v2

@@ 5.4 Results and Statistical Comparisons for Increased Prediction and Sequence Lengths
Table 4: Average MSE and MAE for different prediction lengths and sequence length of 2880 comparing TimeMachine and TTT architectures.
@@
Both shorter and longer sequence lengths affect model performance differently. Shorter sequence lengths (e.g., 2880) provide better accuracy for shorter prediction horizons, with the TTT model achieving an MSE of 0.332 and MAE of 0.356 at horizon 192, outperforming TimeMachine. Longer sequence lengths (e.g., 5760) result in higher errors, particularly for shorter horizons, but TTT remains more resilient, showing improved performance over TimeMachine. For shorter prediction lengths (96 and 192), TTT consistently yields lower MSE and MAE compared to TimeMachine. As prediction lengths grow to 720, both models experience increasing error rates, but TTT maintains a consistent advantage. For instance, at horizon 720, TTT records an MSE of 0.517 compared to TimeMachine’s 0.535. Overall, TTT consistently outperforms TimeMachine across most prediction horizons, particularly for shorter sequences and smaller prediction windows. As the sequence length increases, TTT’s ability to manage long-term dependencies becomes increasingly evident, with models like Conv Stack 5 showing stronger performance at longer horizons.
@@ 5.5 Evaluation
  • Impact of Hidden Layer Architectures: While stacked convolutional architectures, such as Conv Stack 3 and Conv Stack 5, provided incremental improvements, simpler architectures like Conv 3 and Conv 5 also delivered competitive performance. Conv Stack 5 showed a reduction in MSE compared to TimeMachine, at horizon 96, where it achieved an MSE of 0.261 versus TimeMachine’s 0.262. The ModernTCN architecture showed promising results, especially in capturing short-term dependencies effectively.
@@ 6 Conclusion and Future Work
In this work, we improved the state-of-the-art (SOTA) model for time series forecasting by replacing the Mamba modules in the original TimeMachine model with Test-Time Training (TTT) modules, which leverage linear RNNs to capture long-range dependencies. Extensive experiments demonstrated that the TTT architectures—MLP and Linear—performed well, with MLP slightly outperforming Linear. Exploring alternative architectures, particularly Conv Stack 5 and ModernTCN, significantly improved performance at longer prediction horizons, with ModernTCN showing notable efficiency in capturing short-term dependencies. The most significant gains came from increasing sequence and prediction lengths, where our TTT models consistently matched or outperformed the SOTA model, particularly on larger datasets like Electricity, Traffic, and Weather, emphasizing the model’s strength in handling long-range dependencies. While convolutional stacks and ModernTCN showed promise, further improvements could be achieved by refining hidden layer configurations and exploring architectural diversity. Overall, this work demonstrates the potential of TTT modules in long-term forecasting, especially when combined with robust convolutional architectures and applied to larger datasets and longer horizons.
@@ C.2 Effect of Prediction Length
Appendix D Appendix D: Tables
Dataset Channels Time Points Frequencies
Weather 21 52696 10 Minutes
Traffic 862 17544 Hourly
Electricity 321 26304 Hourly
ETTh1 7 17420 Hourly
ETTh2 7 17420 Hourly
ETTm1 7 69680 15 Minutes
ETTm2 7 69680 15 Minutes
Table 6: Details of each dataset
TTT with ModernTCN
ETTh1 ETTh2 ETTm1 ETTm2 Weather Traffic Electricity
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.389 0.402 0.285 0.340 0.322 0.362 0.189 0.273 0.165 0.209 * * * *
192 0.425 0.422 0.359 0.386 0.385 0.397 0.251 0.310 0.265 0.281 * * * *
366 0.460 0.442 0.351 0.388 0.415 0.416 0.309 0.344 0.269 0.293 * * * *
720 0.485 0.475 0.439 0.449 0.490 0.454 0.410 0.403 * * * * * *
Table 7: TTT with ModernTCN results on different datasets. Star symbol indicates that experiment is still running.